# Representation Learning with Parameterised Quantum Circuits for Advancing Speech Emotion Recognition

### v2

@@ Abstract
Speech Emotion Recognition (SER) is a complex and challenging task in human-computer interaction due to the intricate dependencies of features and the overlapping nature of emotional expressions conveyed through speech. Although traditional deep learning methods have shown effectiveness, they often struggle to capture subtle emotional variations and overlapping states. This paper introduces a hybrid classical-quantum framework that integrates Parameterised Quantum Circuits (PQCs) with conventional Convolutional Neural Network (CNN) architectures. By leveraging quantum properties such as superposition and entanglement, the proposed model enhances feature representation and captures complex dependencies more effectively than classical methods. Experimental evaluations conducted on benchmark datasets, including IEMOCAP, RECOLA, and MSP-Improv, demonstrate that the hybrid model achieves higher accuracy in both binary and multi-class emotion classification while significantly reducing the number of trainable parameters. While a few existing studies have explored the feasibility of using Quantum Circuits to reduce model complexity, none have successfully shown how they can enhance accuracy. This study is the first to demonstrate that Quantum Circuits has the potential to improve the accuracy of SER. The findings highlight the promise of QML to transform SER, suggesting a promising direction for future research and practical applications in emotion-aware systems.
@@ I Introduction
Representation learning has emerged as a dominant approach in machine learning, enabling systems to extract meaningful features directly from raw data. This paradigm is particularly significant in the field of Speech Emotion Recognition (SER) as it addresses the complex task of interpreting emotional expressions conveyed through speech. Despite recent advancements, the field still faces challenges in achieving reliable accuracy without complicating the model with high number of parameters. Key factors contributing to these difficulties include the overlapping nature of emotional states, variability among speakers, and the intricate dependencies among various speech features, as noted in the literature [1, 2].
Quantum Machine Learning (QML) provides a transformative perspective on these challenges by leveraging the unique properties of quantum systems, such as superposition and entanglement. Superposition allows quantum systems to process multiple possibilities simultaneously, enabling QML models to analyse overlapping and nuanced emotional cues more effectively and efficiently than traditional methods. Speech signals often contain subtle emotional markers distributed across pitch, rhythm, and tone. While classical systems typically process these markers sequentially or independently, potentially losing important context, superposition allows for the simultaneous consideration of all these features, capturing the complexity of emotional speech more holistically.
Entanglement further enhances QML’s potential in SER by modelling the interdependencies among speech features [3, 4, 5, 6]. Emotional expressions are rarely conveyed through isolated vocal attributes; instead, they arise from intricate relationships between various dimensions of speech, such as tone variations that correlate with tempo or energy shifts. QML utilises entanglements to naturally encode and analyse these dependencies, allowing for a deeper understanding of the nuanced patterns that underpin emotional states.
Parameterised Quantum Circuits (PQCs) are a key innovation in QML, particularly suited to noisy intermediate-scale quantum (NISQ) devices [7]. PQCs are quantum models composed of tunable gates, allowing them to learn complex representations by optimising parameters during training. Unlike traditional deep learning models, which often require a large number of parameters to model high-dimensional spaces, PQCs can encode exponentially larger feature spaces with fewer resources [8]. This capability makes PQCs particularly advantageous for SER, where capturing intricate, high-dimensional emotional dependencies is crucial.
This paper presents a novel framework for SER that integrates PQCs within a representation learning paradigm, as illustrated in Figure 1. By harnessing the unique capabilities of PQCs, our approach aims to address the limitations of classical deep learning methods, including high computational costs and challenges in modelling complex feature dependencies. The proposed method not only improves accuracy but also enhances computational efficiency, presenting a robust solution to one of the most intricate challenges in human-computer interaction.
@@
The key contributions of this study are as follows:
  • A novel SER hybrid classical-quantum framework that integrates PQCs with conventional CNN. This architecture leverages quantum superposition and entanglement to enhance feature representation, addressing the intricate dependencies and overlapping emotional states inherent in speech signals.
  • Compared to purely classical approaches, the proposed model achieves superior performance in both binary and multi-class emotion classification tasks while significantly reducing the number of trainable parameters of the model.
  • Experimental validation across multiple benchmark datasets (IEMOCAP, RECOLA, and MSP-Improv) underscores the efficacy of the hybrid model. The results indicate improved accuracy with smaller number of trainable parameters, setting a strong foundation for future exploration of quantum-enhanced SER applications.
@@ II-A3 Scalability Across Languages and Domains
II-A4 Challenges in Representation Learning for SER: Computational Demands and High-Dimensional Modelling
@@
II-A5 QML addressing the Challenges in Representation Learning for SER
@@ II-B1 Parameterised Quantum Circuits (PQCs)
QML has also shown promise in augmenting traditional neural network structures through Quantum Neural Networks (QNNs). By employing PQCs as neural layers, QNNs can encode quantum states while exploiting quantum advantages for specific tasks. Farhi and Neven pioneered quantum-enhanced learning frameworks for combinatorial optimisation [8], and Schuld and Killoran demonstrated how quantum circuits can efficiently process and generalise high-dimensional data [26], showcasing their suitability for complex learning tasks. While none of the above studies focused on SER, Qu et al. introduced a Quantum Federated Learning (QFL) algorithm for emotion recognition in speech, designed with applications in Internet-on-Vehicles (IoV) technology. Their method employed quantum minimal gated unit RNNs, emphasising privacy and robustness in noisy environments [27]. While this work used PQCs, it did not focus on representation learning.
@@ II-B2 Hybrid Classical-Quantum Architectures
Hybrid quantum-classical architectures have been explored to enhance model expressiveness and training efficiency. These architectures integrate the representational strengths of classical deep learning with the computational power of quantum systems [28, 29]. Hybrid models are particularly suited for tasks where classical models face limitations, as quantum layers can introduce novel features and expand the solution space. Thejha et al. applied a hybrid quantum-classical CNN for speech recognition, demonstrating the potential of hybrid models in audio processing [30]. However, their work did not specifically target SER tasks. Similarly, Esposito et al. employed QNNs to process MFCC features for audio classification, passing the learnt representations through an RNN for COVID-19 cough detection [31]. While this study utilised hybrid architectures and PQCs, it did not focus on representation learning for SER. Similarly, Norval and Wang proposed a Quantum AI-based approach for SER, utilising features like MFCC, pitch, and energy. Their system employed quantum circuits for encoding and variational quantum algorithms for classification, achieving an accuracy of 30% on a custom dataset [32]. However, their approach did not investigate PQCs or representation learning for SER.
An area of significant interest is the application of hybrid architectures in CNNs, leading to the development of quantum CNNs (QCNNs). QCNNs effectively process structured data by utilising quantum circuits to perform convolution-like operations, as demonstrated by Cong et al. [33]. Further advancements have shown the applicability of QCNNs to classical data domains, including SER. Thejha et al. extended QCNNs to SER tasks, achieving performance comparable to classical CNNs [30].
@@ III-A Preamble
Quantum Measurement is the process of extracting classical information from quantum states by projecting them onto specific basis states [25]. In QNNs, measurements occur at the end of the quantum circuit to produce outputs such as classifications, probabilities, or projections. The choice of measurement basis significantly influences the interpretability and effectiveness of the results. In this study, we employ the following quantum measurement methods: PauliZ, PauliX, Probability, and Z.
PauliZ Measurement – corresponds to measuring a qubit in the computational (standard) basis, $\{|0\rangle,|1\rangle\}$. This measurement projects the qubit state onto the eigenstates of the Pauli Z operator: $|0\rangle(eigenvalue+1)$ and $|1\rangle(eigenvalue-1)$. The mathematical representation of the PauliZ measurement is as follows.
$Z=\begin{bmatrix}1&0\\ 0&-1\end{bmatrix}$
Z Measurement – is shorthand for a measurement in the Pauli Z basis, projecting the state onto the $|0\rangle$ and $|1\rangle$ eigenstates of the Z operator. It essentially measures whether the qubit state is $|0\rangle$ or $|1\rangle$, directly giving the classical bit value. The difference between PauliZ and Z measurements are how the results are interpreted or used: as eigenvalues of an operator (+1, -1) or as classical binary outcomes (0, 1).
@@ III-B Model Architecture
Figure 3 illustrates the quantum circuit of a strongly entangling layer with four qubits. This configuration exemplifies how entanglement is incorporated into the quantum layer to improve the model’s ability to learn nuanced emotional features from speech data.
@@ IV-A Datasets
We use three publicly available datasets commonly utilised in the field of SER: IEMOCAP [38], MSP-Improv [40], and RECOLA [39].
@@ IV-A1 IEMOCAP
is a widely used dataset containing 12 hours of acted multi-speaker audio-visual data, with dyadic sessions that include both scripted and improvised scenarios. It provides categorical emotion labels (happiness, sadness, anger, and neutrality) and dimensional labels (valence, dominance, and activation). In this study, we focus on the audio modality from the improvised scenarios and specifically use the valence annotations for analysis. The valance value of this dataset annotation ranges from 1 to 5.
@@ IV-A2 MSP-Improv
is a widely recognised audio-visual dataset frequently employed in multi modal SER research. Initially developed for a study on audio-visual emotional perception, it has since been utilised in numerous SER studies [45, 46, 47, 48], making it a suitable choice for our research. The dataset was recorded in a controlled setting and features 20 pre-determined scripts designed to represent key emotions, including happiness, sadness, anger, and neutrality.
@@ IV-A3 RECOLA
is a multi-modal dataset designed for emotion recognition, containing recordings of naturalistic interactions between pairs of participants. It includes audiovisual and physiological data, annotated with both categorical emotions and dimensional labels, including valence. In this study, we use the valence annotations of RECOLA in our experiments.
@@ V-C Summary of the Results
A summary of the results obtained by the experiments in this study is tabulated in the Table III. The results demonstrate the effectiveness of the hybrid classical-quantum model for SER tasks, highlighting superior performance and reduced complexity compared to classical models. For binary classification on the IEMOCAP dataset, the hybrid model achieved a UAR of 64.68%, outperforming the classical model’s 61.36%, with $\sim$50fewer trainable parameters.
@@ VII Conclusion
The experimental results across three widely used benchmark datasets – IEMOCAP, RECOLA, and MSP-Improv – consistently showcase the superior performance of the hybrid classical-quantum model compared to its purely classical counterpart. Notably, the hybrid models achieve enhanced accuracy in both binary and multi-class emotion classification tasks, while simultaneously reducing the number of trainable parameters by almost half. This reduction in model complexity translates to lower computational overhead and improved energy efficiency, rendering these quantum-enhanced methods more suitable for practical real-world applications.
Furthermore, the study reveals crucial insights into optimal configurations for hybrid classical-quantum models in SER. The experiments suggest that specific quantum embeddings, such as Angle and Amplitude embedding, coupled with strongly entangling or random quantum circuit layers, and a summation of Pauli Z measurements consistently contribute to superior performance. The absence of weight decay in these optimal configurations implies that L2 regularisation might not be necessary for such hybrid models. These findings provide valuable guidance for future research and development in quantum-enhanced SER.

### v3

@@ Abstract
Quantum machine learning (QML) offers a promising avenue for advancing representation learning in complex signal domains. In this study, we investigate the use of parameterised quantum circuits (PQCs) for speech emotion recognition (SER)—a challenging task due to the subtle temporal variations and overlapping affective states in vocal signals. We propose a hybrid quantum-classical architecture that integrates PQCs into a conventional convolutional neural network (CNN), leveraging quantum properties such as superposition and entanglement to enrich emotional feature representations. Experimental evaluations on three benchmark datasets—IEMOCAP, RECOLA, and MSP-IMPROV—demonstrate that our hybrid model achieves improved classification performance relative to a purely classical CNN baseline, with over 50% reduction in trainable parameters. This work provides early evidence of the potential for QML to enhance emotion recognition and lays the foundation for future quantum-enabled affective computing systems.
@@ I Introduction
Representation learning has become a foundational approach in machine learning, enabling systems to extract meaningful patterns directly from raw data. In the field of Speech Emotion Recognition (SER), this paradigm is particularly valuable for interpreting the rich and often ambiguous emotional signals embedded in vocal expressions. Despite recent progress, SER continues to face significant challenges in achieving high accuracy without resorting to increasingly complex models with a large number of trainable parameters. Key obstacles include the overlapping nature of emotional states, speaker variability, and the intricate interdependencies among acoustic features such as pitch, energy, and rhythm [1, 2, 3].
Quantum Machine Learning (QML) offers a compelling alternative by exploiting the unique computational characteristics of quantum systems: most notably, superposition and entanglement. Superposition enables quantum models to process multiple states simultaneously, allowing for a more holistic analysis of the nuanced and concurrent variations inherent in emotional speech. Rather than evaluating features in isolation or sequentially, as is typical in classical systems, quantum models can assess complex patterns across all dimensions in parallel.
Entanglement further enhances this potential by capturing the dependencies between speech features that classical models often overlook or treat independently [4, 5, 6, 7]. In practice, emotional states are rarely expressed through isolated vocal characteristics; instead, they emerge from subtle relationships between multiple attributes, such as correlations between tone, tempo, and energy. QML systems inherently encode and analyse these relationships, leading to richer and more robust representations of affective information.
Parameterised Quantum Circuits (PQCs) are among the most promising constructs in QML. These circuits consist of tunable quantum gates, which can be trained to optimise complex representations. Unlike conventional deep neural networks—which often rely on large numbers of parameters to model high-dimensional spaces—PQCs can encode exponentially large feature spaces using comparatively fewer resources [8]. This makes them especially suitable for SER, where modelling emotional intricacies efficiently is critical.
The emergence of Noisy Intermediate-Scale Quantum (NISQ) devices, such as quantum processors with limited qubit counts and moderate noise resilience, has made it feasible to explore hybrid quantum-classical learning strategies. While current quantum hardware is not yet mature for large-scale deployment, simulations of PQC-based models offer valuable insights into their performance and architectural viability under NISQ constraints.
In this study, we propose a novel hybrid architecture for speech emotion recognition that integrates PQCs within a classical convolutional neural network framework, as illustrated in Figure 1. By combining the expressive power of quantum circuits with the proven feature extraction capabilities of CNNs, our approach addresses key limitations of classical models in terms of both accuracy and computational efficiency. Experiments conducted on three benchmark datasets—IEMOCAP, RECOLA, and MSP-IMPROV—demonstrate that the hybrid model achieves improved classification performance compared to a classical CNN baseline, while reducing the number of trainable parameters by over 50%.
This work contributes to the growing field of quantum-enhanced affective computing by demonstrating that PQCs can be effectively applied to emotionally rich, temporally structured data. It also provides empirical insights into how quantum circuit design, particularly expressibility and entanglement, affects model performance, laying the groundwork for future implementation on real quantum hardware.
@@ II-A2 Enhanced Representational Power
Further advancing classical methods, researchers have also explored complex architectures to address multiple SER challenges simultaneously. For instance, Daneshfar and Kabudian proposed a system using a hierarchical multi-layer sparse auto-encoder combined with an Extreme Learning Machine (ELM) [2]. Their approach utilised a rich set of spectral and spectro-temporal features and notably introduced a new adaptive weighting method specifically to counteract the problem of data imbalance, a persistent challenge in SER datasets. This highlights a direction in classical SER research focused on creating sophisticated, multi-stage processing pipelines to extract discriminative features and build robust classifiers.
@@ II-A4 Robustness to Limited Labelled Data
II-A5 Challenges in Representation Learning for SER: Computational Demands and High-Dimensional Modelling
@@
II-A6 QML addressing the Challenges in Representation Learning for SER
@@ II-B1 Parameterised Quantum Circuits (PQCs)
QML has also shown promise in augmenting traditional neural network structures through Quantum Neural Networks (QNNs). By employing PQCs as neural layers, QNNs can encode quantum states while exploiting quantum advantages for specific tasks. Farhi and Neven pioneered quantum-enhanced learning frameworks for combinatorial optimisation [8], and Schuld and Killoran demonstrated how quantum circuits can efficiently process and generalise high-dimensional data [27], showcasing their suitability for complex learning tasks. While none of the above studies focused on SER, Qu et al. introduced a Quantum Federated Learning (QFL) algorithm for emotion recognition in speech. Their method emphasised privacy and robustness in noisy environments by employing quantum minimal gated unit RNNs [28]. However, a key shortcoming of this work is that it did not focus on representation learning. By not leveraging the feature representation capabilities of PQCs, the model’s ability to capture the complex, high-dimensional dependencies inherent in emotional speech data remained unexplored, limiting its potential effectiveness for the core SER task.
@@ II-B2 Hybrid Classical-Quantum Architectures
Hybrid quantum-classical architectures have been explored to enhance model expressiveness and training efficiency. These architectures integrate the representational strengths of classical deep learning with the computational power of quantum systems [29, 30]. Hybrid models are particularly suited for tasks where classical models face limitations, as quantum layers can introduce novel features and expand the solution space. Other explorations into hybrid quantum-classical architectures have shown promise but have not yet unlocked a quantum advantage for SER. For instance, Thejha et al. applied a hybrid quantum-classical CNN to general speech recognition, achieving performance merely comparable to classical CNNs without demonstrating a clear benefit [31]. Likewise, work by Esposito et al. on audio classification for cough detection [32] also failed to investigate representation learning for SER. These studies, while foundational, stopped short of demonstrating that QML could either outperform classical methods or effectively model the nuanced feature representations required for robust emotion recognition. Similarly, Norval and Wang proposed a Quantum AI-based approach for SER, but its practical viability was limited, achieving a low accuracy of only 30% on a custom dataset [33]. Their approach, which relied on basic quantum encoding and variational algorithms without exploring PQCs or representation learning, underscores the significant performance challenges that early QML-SER models faced.
An area of significant interest is the application of hybrid architectures in CNNs, leading to the development of quantum CNNs (QCNNs). QCNNs effectively process structured data by utilising quantum circuits to perform convolution-like operations, as demonstrated by Cong et al. [34].
@@ III-A Preamble
Quantum Measurement is the process of extracting classical information from quantum states by projecting them onto specific basis states [25]. In QNNs, measurements occur at the end of the quantum circuit to produce outputs such as classifications, probabilities, or projections. The choice of measurement basis significantly influences the interpretability and effectiveness of the results. In this study, we employ the following quantum measurement methods: PauliZ, Z, PauliX, and Probability. The choice of measurement is critical as it determines the nature of the classical data passed to the subsequent layers of our model.
PauliZ Measurement – calculates the expectation value of the Pauli Z operator, which is the average value of its eigenvalues (+1 for state $|0\rangle$ and -1 for state $|1\rangle$). This results in a continuous output value in the range $[-1,+1]$. This continuous output can be interpreted as a probabilistic score or a “soft” classification from the quantum circuit. For an SER task, this allows the model to express the degree to which an emotional feature is present, rather than making a hard decision. This nuanced output is then fed into the classical neural network for further processing to reach a final classification [42].
Z Measurement – is a measurement in the computational basis, which forces the qubit into a definite state of either $|0\rangle$ or $|1\rangle$. This process, known as state collapse, yields a classical binary outcome (a 0 or a 1)![43]. In the context of binary SER classification (e.g., ’High’ vs. ’Low’ valence), this provides a direct, decisive classification from the quantum layer. This binary output can then be used by the subsequent classical layers, potentially simplifying the final classification decision.
The fundamental difference lies in their output: PauliZ provides a continuous, probabilistic value, while Z provides a discrete, binary outcome. This distinction directly affects how the SER model interprets the quantum-processed features, with PauliZ offering a more nuanced, probabilistic representation and Z providing a more direct, binary classification.
@@ III-B Model Architecture
We acknowledge the limitations of this approach. The baseline CNN model does not incorporate more complex components like LSTM layers for temporal modeling or attention mechanisms, which have been shown to enhance SER performance by enabling the model to dynamically focus on the most emotionally salient segments of speech [46]. This choice represents a critical trade-off: while a more sophisticated classical architecture could potentially yield higher overall accuracy, it would also introduce confounding variables, making it more difficult to isolate and interpret the specific impact of the quantum components. Our primary goal in this study is to establish a clear and foundational understanding of the quantum layer’s capabilities. By benchmarking against a simpler, well-understood classical model, we can more directly attribute performance gains to the quantum representation learning, thereby balancing the pursuit of absolute performance with the need for clear, interpretable results.
@@
Figure 3 illustrates the quantum circuit of a strongly entangling layer with four qubits. This configuration exemplifies how entanglement is incorporated into the quantum layer to improve the model’s ability to learn nuanced emotional features from speech data. The design of this four-qubit circuit, which uses sequential controlled-NOT (CNOT) gates to link the qubits, systematically builds up correlations between them. This robust entanglement is critical for the model’s ability to represent the complex, interdependent relationships among the various acoustic features that characterise emotional speech.
@@ IV-A Datasets
To robustly evaluate the proposed hybrid classical-quantum model, we selected three benchmark datasets commonly used in the SER field: IEMOCAP [39], MSP-Improv [41], and RECOLA [40]
@@ IV-A1 IEMOCAP
is a widely used dataset containing 12 hours of audio-visual data from dyadic interactions between actors. The sessions include both scripted and improvised scenarios designed to elicit specific emotions. While the emotions are acted, the improvisational nature of many interactions introduces a degree of spontaneity not found in purely scripted datasets. By including IEMOCAP, we test our model’s performance on more archetypal and clearly expressed emotions, which serves as a crucial baseline for SER tasks. In this study, we focus on the audio from the improvised sessions and use the valence annotations, which range from 1 to 5, to evaluate the model’s ability to capture nuanced emotional states.
@@ IV-A2 MSP-Improv
is a valuable resource designed to bridge the gap between acted and naturalistic data. It features actors improvising scenarios that are specifically designed to elicit genuine emotional responses in a controlled setting. This methodology aims to produce more realistic emotional expressions than simple script-reading, providing a unique test case for our model. The dataset contains 20 predetermined scripts, allowing for the study of key emotions like happiness, sadness, anger, and neutrality. Using MSP-Improv allows us to evaluate our model’s effectiveness on lexically controlled but emotionally varied speech.
@@ IV-A3 RECOLA
consists of recordings of spontaneous, naturalistic interactions between participants collaborating on a task remotely. The emotional expressions in RECOLA are not acted or elicited, but arise naturally from the interaction, making them more subtle and representative of real-world conversational speech. This presents a significant challenge for any SER model. By evaluating on RECOLA’s valence annotations, we specifically test our model’s capacity to recognise emotions in spontaneous and unscripted contexts, which is a key objective for advancing the field of SER.
@@ IV-C Model Configuration
The models were implemented using the PyTorch framework, with the quantum components simulated using the PennyLane library. All training and evaluation experiments were conducted on a high-performance computing node equipped with an NVIDIA A100 40GB GPU, a 60-core CPU, and 120GB of RAM. Given the extensive hyperparameter space, the grid search was computationally intensive, requiring significant runtime to systematically evaluate all possible configurations and identify the optimal model for each experiment.
@@ V-C Summary of the Results
A summary of the results obtained by the experiments in this study is tabulated in the Table III. The results demonstrate the effectiveness of the hybrid classical-quantum model for SER tasks, highlighting superior performance and reduced complexity compared to classical models. For binary classification on the IEMOCAP dataset, the hybrid model achieved a UAR of 64.68%, outperforming the classical model’s 61.36%, with 50.34% fewer trainable parameters.
@@ VII Conclusion
The experimental results across three widely used benchmark datasets – IEMOCAP, RECOLA, and MSP-Improv – consistently showcase the superior performance of the hybrid classical-quantum model compared to its purely classical counterpart. Notably, the hybrid models achieve enhanced accuracy in both binary and multi-class emotion classification tasks, while simultaneously reducing the number of trainable parameters by 50.34%. This reduction in model complexity translates to lower computational overhead and improved energy efficiency, rendering these quantum-enhanced methods more suitable for practical real-world applications.
Furthermore, the study reveals crucial insights into optimal configurations for hybrid classical-quantum models in SER. The experiments suggest that specific quantum embeddings, such as Angle and Amplitude embedding, coupled with strongly entangling or random quantum circuit layers, and a summation of Pauli Z measurements consistently contribute to superior performance. The absence of weight decay in these optimal configurations implies that L2 regularisation might not be necessary for such hybrid models. Notably, the optimal configurations consistently emerge with zero weight decay. This directly supports the observation from Section VI-A, which indicates that the quantum layers inherently provide sufficient regularisation. This behaviour suggests that classical techniques like L2 regularisation may be redundant, potentially due to the intrinsic properties and constraints of the quantum Hilbert space, simplifying the training process. These findings provide valuable guidance for future research and development in quantum-enhanced SER.